OpenStudy (anonymous):

I know this is SO silly & this sounds ridiculous to answer but can someone PLEASE help me with this story problem? Reflect upon, analyze, and critique students statements. Steph stated to Alya that her age is the solution to the equation “ the square root of x plus four is equal to four!” To which Alya replied, "No way! That means that you were born the instant you stated your equation!" Do you agree with Alya, with Steph, or with neither one of them? Justify your response.

OpenStudy (phi):

Nobody has conversations like that! But let's go with it... There is ambiguity in what Steph is saying. When she says the square root of x + 4 = 4, does she mean \[ \sqrt{x+4}=4\] or \[ \sqrt{x}+4=4\] If you assume the first, then we see x=12 is a valid answer. If we assume the latter, then x=0. So it appears that Alya assumed the second interpretation. I would say neither is wrong, but rather this is a failure to communicate.

OpenStudy (anonymous):

thanks so much